### v1

# When Does Divide and Conquer Work for Long Context LLM? A Noise Decomposition Framework

Abstract
We investigate the challenge of applying Large Language Models (LLMs) to long texts. We propose a theoretical framework that distinguishes the failure modes of long context tasks into three categories: cross-chunk dependence (task noise), confusion that grows with context size (model noise), and the imperfect integration of partial results (aggregator noise). Under this view, we analyze when it is effective to use multi-agent chunking, i.e., dividing a length sequence into smaller chunks and aggregating the processed results of each chunk. Our experiments on tasks such as retrieval, question answering, and summarization confirm both the theoretical analysis and the conditions that favor multi-agent chunking. By exploring superlinear model noise growth with input length, we also explain why, for large inputs, a weaker model configured with chunk-based processing can surpass a more advanced model like GPT4o applied in a single shot. Overall, we present a principled understanding framework and our results highlight a direct pathway to handling long contexts in LLMs with carefully managed chunking and aggregator strategies [^1] The code implementation will be available at https://github.com/NehzUx/DivideConquerAgents.
1 Introduction
Large Language Models (LLMs) have drawn significant interest from both industry and academia, thanks to their ability to handle tasks ranging from open-ended question answering to complex reasoning. As these models grow in capacity, there is an increasing demand to apply them to extended texts that may span hundreds of thousands of tokens. In principle, self-attention architectures are powerful, but their reliance on quadratic operations in sequence length can make long-context tasks computationally expensive Tay et al. (2022). Moreover, even if a model technically can process long contexts, studies have reported the quality decline of output once the input surpasses a certain length (Hsieh et al., 2024). This limitation has been attributed to phenomena such as the “lost in the middle” effect, where the model forgets or mishandles portions of the input.
Existing research has explored ways to ease these difficulties. One type of approach has focused on modifying the transformer architecture to reduce memory footprint and compute, often by, for example, shaping the attention pattern through blockwise or window-based strategies (Qiu et al., 2020; Beltagy et al., 2020), low-rank approximations (Wang et al., 2020) or routing-based approaches (Kitaev et al., 2020), seeking to prolong the feasible context length without harming accuracy too severely. Although these technical improvements often extend the maximum input size, they do not guarantee stable performance when that size becomes very large.
A more functional approach divides a large input into chunks, processes each chunk with one or more worker models, and then aggregates the outputs. Retrieval-augmented pipelines are a popular example of this idea (Lewis et al., 2020; Fan et al., 2024; Wang et al., 2024b), but they often rely on ad hoc rules for aggregating results. Their effectiveness hinges on how well global dependencies are preserved. If the aggregation step is weak, cross-chunk reasoning can be lost, which leads to suboptimal outcomes. While these chunk-based methods can lower confusion at the worker level, the completeness of the final result is not guaranteed.
In this paper, we propose a divide and conquer framework Qian et al. (2024b); Zhang et al. (2024b); Zhou et al. (2024) that offers a more systematic view at these issues. we analyze these challenges by establishing a theoretical framework that tracks three main sources of error in long-context tasks. First, task noise arises from cross-chunk dependencies that cannot be handled by processing each segment alone. Second, model noise arises from degraded performance with increasing input length. Third, aggregator noise emerges when partial results are combined incorrectly, even if each chunk was handled well. We show how the balance of these factors determines when chunk-based division is advantageous, and how the design of the aggregator stage plays an important role. In particular, when a task demands only mild cross-chunk reasoning, chunking can reduce model confusion with little downside. However, when cross-chunk synergy is large, only a more advanced aggregator can preserve important global connections.
We also show that model noise can worsen at a rate that outstrips the benefit of seeing the entire input at once. Beyond a certain length threshold, splitting the text can yield better performance, even if the individual worker models are weaker than a single large model. We attribute this to a superlinear increase in model error growth with input length. Our experiments on retrieval, QA, summarization, and other tasks confirm that this superlinear effect surfaces in practice, and demonstrate how a planner can arrange prompts to keep aggregator noise manageable. Across these long context tasks, weak models handling chunks often outperform a single-shot strong model (e.g. GPT4o).
Our contributions are as follows:
  • We present a theoretical framework for modeling error terms in long-context processing. This framework helps to explain when multi-agent chunking is advantageous and when it is not.
  • We provide empirical evidence of superlinear performance degradation when input length is large. We then demonstrate that splitting into smaller pieces can help mitigate these effects, except in the case where cross-chunk synergy is very high.
  • We show that carefully designed prompts for both worker agents and the aggregator can stably improve final performance. This allows smaller models to surpass more advanced models on certain tasks when the input length is large.
These findings broaden our understanding of how to tackle large inputs with LLMs. By breaking down the main sources of error, our framework clarifies why chunking can help, how it can fail, and how a planner can structure prompts so that partial outputs lead to correct final answers. Through both theoretical arguments and experiments, we reveal practical guidelines for handling lengthy contexts and show that well-planned division of labor can be a strong alternative to single-shot processing with a massive context window.
2 Related works
2.1 Long context LLM
Developing long context LLMs has seen increasing popularity in applications such as long document understanding. The transformer architecture plays a dominating role in language modeling and, due to the attention mechanism, it is associated with quadratic computational and memory complexity against sequence length(Tay et al., 2022). Therefore, it is expensive to train and serve long context LLMs in practice. To mitigate this, researchers have explored various efficient transformer architectures, such as block wise attention(Qiu et al., 2020), window attention(Beltagy et al., 2020), low-rank approximated attention(Wang et al., 2020), routing attention(Kitaev et al., 2020), etc. (Dao et al., 2022) proposed flash attention and (Liu et al., 2023) designed ring attention, both contributed to efficient training of long context LLMs. The recent progress on transformer alternatives also show promises of overcoming the computational complexity of attention mechanisms, including state-space models, e.g, Mamba(Gu and Dao, 2024) and convolutional-based approach, Hyena(Poli et al., 2023). On the other hand, given existing LLMs trained on short context, many research works have been presented on length extrapolation based on positional encoding(Chen et al., 2023; Peng et al., 2024; Jin et al., 2024), and retrieval-augmented generation(Lewis et al., 2020; Wang et al., 2024b), where additional design complexity is be added on positional encoding adaptation and retriever.
2.2 Multi-Agents
Recent research has explored the potential of multi-agent systems utilizing large language models (LLMs) for complex task-solving. In the realm of task decomposition and planning, Qian et al. (2023) proposed a framework where multiple LLM-based agents collaboratively break down tasks into subtasks and develop execution plans. Similarly, Haji et al. (2024) introduced a hierarchical mechanism where tree of thoughts reasoners are combined with thought verifiers. Zhao et al. (2024b) breaks down and process long texts using multiple agents, enabling efficient handling of 128K token contexts through task decomposition.
Beyond task breakdown and planning, recent work focused on communication mechanisms within multiple agents. Several studies (Guo et al., 2024) have investigated various approaches to enable effective information exchange between agents. Common communication paradigms such as debate is commonly used (Du et al., 2024; Chan et al., 2024; Liang et al., 2023). Alternatively, Qian et al. (2024a) adapts a cooperative paradigm where agents work together towards a common goal. Many multi-agent works use a layered structure. Wang et al. (2024a) developed a multi-layered framework that separates LLMs into proposers and aggregators. Liang et al. (2023) encourages more divergent thinking though layered debate among debtors. However, a centralized sharing structure has also been explored by Hong et al. (2023).
2.3 Divide and conquer
Divide-and-conquer is a classic computational strategy that has recently been integrated with LLM-based multi-agent systems to handle long-context processing. Recent works such as LC-Boost Qian et al. (2024b), Chain-of-Agents (CoA) Zhang et al. (2024b), LongAgent Zhao et al. (2024a), and LLM×$\times$×MapReduce Zhou et al. (2024) , adopt this paradigm by splitting long inputs into smaller chunks processed by worker agents and then merging their outputs through a manager agent. These methods improve efficiency and enable long-text reasoning without requiring large context models.
However, existing approaches lack a formal theoretical framework to analyze the interaction between task complexity, model noise, and aggregation errors, making it difficult to optimize chunking strategies. They also struggle with understanding how cross-chunk dependencies impact performance, often leading to loss of contextual coherence when aggregating local outputs. Additionally, agent coordination mechanisms in prior work are heuristic-driven, lacking a principled way to balance worker and manager responsibilities. This work addresses these gaps by providing a rigorous modeling framework for analyzing noise propagation in Divide-and-conquer architectures, offering a systematic approach to optimizing multi-agent collaboration for long-context tasks.
3 Theoretical modeling framework
3.1 Problem statement and setting
We consider a long input $x$ of length $T$ tokens, denoted $x\in\mathcal{X}$, where $\mathcal{X}$ is the input space. The input $x$ can be split into $n$ contiguous chunks: $x\;=\;(x_{1},x_{2},\dots,x_{n}),\quad\text{with}\quad\sum_{i=1}^{n}\|x_{i}\|\;=\;T.$ A task asks for a mapping function $f:\mathcal{X}\to\mathcal{Y}$ where $\mathcal{Y}$ is the answer space. The target task function $y=f^{*}(x)$ is the correct or desired output for the entire input, for example returning the correct answer to a user query or extracting complex reasoning from a text of length $T$.
Single-shot long context agent inference.
We feed the entire $x$ (length $T$) into one agent that has context length equal to or longer than $T$. It computes $f_{\mathrm{single}}(x)\;=\;f^{*}(x)\;+\;\phi_{\mathrm{model}}(x),$ where $\phi_{\mathrm{model}}(x):\mathcal{X}\to\mathcal{Y}$ is a small-phi noise term capturing model errors (for example attention constraints or distribution mismatch). We denote $\phi_{\mathrm{model}}(x)=\phi(\|x\|)$, with $\phi(\cdot)$ non-decreasing. In practice, performance can degrade with $T$ even if $T$ is within the nominal context window.
Divide and conquer agents (D&C).
We split $x$ into $n$ subproblems $(x_{1},\dots,x_{n})$. Each subproblem is handled by a worker agent model $f_{i}(\cdot)$: $f_{i}(x_{i})\;=\;f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i}).$ For each chunk $x_{i}$, the worker agent noise is $\phi_{\mathrm{model}}(x_{i})$. If $\|x_{i}\|=\tfrac{T}{n}$ and $\phi(\cdot)$ grows faster than linearly (shown later), then single-shot usage over $T$ can be significantly noisier than chunk usage over $\tfrac{T}{n}$. Next, an aggregator (manager agent) $h(\cdot)$ merges these sub-results:
$f_{\mathrm{D\&C}}(x;n)$ $=h\bigl{(}f_{1}(x_{1}),\dots,f_{n}(x_{n})\bigr{)}=\prod_{i=1}^{n}f_{i}(x_{i})+\phi_{\mathrm{agg}}(x_{1},\dots,x_{n}).$
If the aggregator $h$ is not fully capable of reconstructing global interactions, an additional noise $\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$ appears in the final output.
3.2 Task noise, model noise, and aggregator noise
We posit that many tasks can be partially factorized across chunks plus a noise term. Let
$g(x_{1},\dots,x_{n})\;=\;\underbrace{\prod_{i=1}^{n}g^{*}(x_{i})}_{\text{factorized component}}\;+\;\underbrace{\phi_{\mathrm{task}}(x_{1},\dots,x_{n})}_{\begin{subarray}{c}\text{cross-chunk dependency}\\ \text{or synergy}\end{subarray}},$
where $g^{*}(x_{i})$ is the local contribution from chunk $x_{i}$, and $\phi_{\mathrm{task}}(\cdot)$ measures how much $g$ departs from being a product of local chunk functions. If $\phi_{\mathrm{task}}\approx 0$, the task is nearly chunk-independent. Otherwise, large $\phi_{\mathrm{task}}$ indicates significant cross-chunk synergy.
When $\|\phi_{\mathrm{task}}\|\leq\epsilon$, a well-aligned aggregator $h$ can keep $\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$ on the order of $\epsilon$. Conversely, if $\phi_{\mathrm{task}}$ is large and the aggregator does not reprocess the entire $x$, then $\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$ may scale with $\|\phi_{\mathrm{task}}\|$.
Putting it all together. Summarizing the contributions of each term in D&C:
$f_{\mathrm{D\&C}}(x;n)$ $=\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}+\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$
$=\prod_{i=1}^{n}f_{i}^{*}(x_{i})+\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})+\Bigl{[}\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}-\prod_{i=1}^{n}f_{i}^{*}(x_{i})\Bigl{]}$
For simplicity, we then define $\Phi_{\mathrm{model}}:=\Bigl{[}\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}-\prod_{i=1}^{n}f_{i}^{*}(x_{i})\Bigl{]},\Phi_{\mathrm{agg}}:=\phi_{\mathrm{agg}}(x_{1},\dots,x_{n}),\Phi_{\mathrm{task}}:=-\phi_{\mathrm{task}}(x_{1},\dots,x_{n}).$ $\Phi_{\mathrm{model}}$ arises from model noise (and is a polynomial in the terms $\phi_{\mathrm{model}}(x_{i})$). $\Phi_{\mathrm{agg}}$ arises from aggregator limitations. $\Phi_{\mathrm{task}}$ is the negative of the task factorization mismatch.
Hence, we have $f_{\mathrm{D\&C}}(x;n)=\prod_{i=1}^{n}f_{i}^{*}(x_{i})+\Phi_{\mathrm{model}}+\Phi_{\mathrm{agg}}.$ We can compare it to the target $g(x)=\prod_{i=1}^{n}g^{*}(x_{i})+\phi_{\mathrm{task}}(x_{1},\dots,x_{n}).$ Then the approximation error is
$f_{\mathrm{D\&C}}(x;n)-g(x)$ $=\underbrace{\Bigl{[}\prod_{i=1}^{n}f_{i}^{*}(x_{i})-\prod_{i=1}^{n}g^{*}(x_{i})\Bigr{]}}_{\Delta^{*}(x)}\quad+\bigl{(}\,\Phi_{\mathrm{model}}\;+\;\Phi_{\mathrm{agg}}\;+\;\Phi_{\mathrm{task}}\,\bigr{)}.$ (1)
$\Delta^{*}(x)$ is the difference between the product of ideal worker outputs and the product of true chunk-level targets. If each $f_{i}^{*}(x_{i})$ matches $g^{*}(x_{i})$ well, the main discrepancy is $\Phi_{\mathrm{model}}+\Phi_{\mathrm{agg}}+\Phi_{\mathrm{task}}$. By improving worker models (to shrink $\phi_{\mathrm{model}}$) and using better aggregation (to reduce $\phi_{\mathrm{agg}}$), D&C can lower overall error. However, a large $\phi_{\mathrm{task}}$ still requires effective cross-chunk merging.
3.3 Why weaker models can outperform stronger ones
We first make precise the statement that $\phi_{\mathrm{model}}(x)$ grows faster than linearly in the input length, eliminating linear or sublinear growth cases:
Assumption 3.1 (Superlinear model noise).
Denote $\phi_{\mathrm{model}}(x)$ the model noise for an input of length $L=\|x\|$. For sufficiently large $L$, there exist constants $c>0$ and $\alpha>1$ such that $\phi_{\mathrm{model}}(x)\geq cL^{\alpha}.$
Justification by elimination. Please refer to Appendix A for details.
Theorem 3.2 (When D&C agents win).
Consider an input $x$ of length $T$. Suppose that the single-shot noise $\phi_{\mathrm{model}}(T)$ grows superlinearly, and that $\Delta^{*}(x)$, the aggregator noise $\Phi_{\mathrm{agg}}$, and the cross-chunk mismatch $\Phi_{\mathrm{task}}$ all remain bounded when the input is divided into $n$ chunks of length at most $L_{0}$. Then, for sufficiently large $T$, one can choose $n\approx T/L_{0}$ to ensure: $|f_{\mathrm{D\&C}}(x;n)-g(x)|<\phi_{\mathrm{model}}(T).$
Proof. Please refer to Appendix B for details.
3.4 Three regimes of noises
In many practical pipelines, $\Phi_{\mathrm{agg}}$ can be kept modest through careful prompt design by a planner agent. Thus, the major interplay often arises between $\Phi_{\mathrm{model}}$ (how confused the model is on each chunk) and $\Phi_{\mathrm{task}}$ (how strongly the chunks depend on each other). We characterize three regimes.
  • Case 1: Both $\Phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$ are negligible. Here, each chunk is easy for the model, and the task factorization mismatch is minor. Splitting into chunks or using a single shot yields similar performance. Even if there is a slight aggregator noise $\Phi_{\mathrm{agg}}$, it stays negligible compared to the small $\Phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$.
  • Case 2: $\Phi_{\mathrm{task}}$ dominates. The input’s synergy is extensive, yet each individual chunk is straightforward. A naive aggregator might fail to merge the global context effectively, causing $\Phi_{\mathrm{agg}}$ to blow up unless prompted carefully. In this regime, single-shot usage could still do well if the model can handle long context, but D&C requires a robust aggregator for small $\Phi_{\mathrm{agg}}$.
  • Case 3: $\Phi_{\mathrm{model}}$ dominates. The task is nearly chunk-independent, but each chunk is intrinsically confusing to the model when taken in a single large context. Splitting the input reduces the per-chunk confusion, lowering $\Phi_{\mathrm{model}}$ significantly. Even if $\Phi_{\mathrm{agg}}$ grows somewhat, it is outweighed by the model noise improvement. This is precisely the domain where Thm 3.2 shows that D&C can outperform a single-shot approach.
These three cases illustrate that neither $\Phi_{\mathrm{task}}$ nor $\Phi_{\mathrm{model}}$ alone determines whether multi-agent chunking or single-shot usage is more suitable. Their relative magnitudes, together with $\Delta^{*}(x)$ and the aggregator’s effectiveness ($\Phi_{\mathrm{agg}}$), ultimately decide which approach is advantageous.
4 A simple implementation of the framework
We develop a minimal three-part system as illustrated in Figure 1: a planner to manage how the original query is segmented and how prompts are arranged, multiple worker agents that each handle a different sub-chunk, and a single manager agent that merges all sub-results. Below we briefly introduce the worker and manager roles, followed by a more detailed description of the planner.
Figure 1: A simple implementation of the divide and conquer framework. The Math task example (right panel) illustrates the planner’s critical role by translating instructions of returning 2nd smallest number into returning the two smallest numbers per chunk.
Worker agent: We split the input sequence into chunks of approximately equal length, and each worker agent is assigned a single chunk. Each worker focuses solely on its segment, without managing cross-segment dependencies. Although we use identical models for the workers, one can easily extend the architecture to mix different worker models as needed.
Manager agent: After the worker agents produce partial outputs, the manager agent merges them into one final result. In our baseline, this manager agent is the same model type as the workers, but one may employ a more specialized manager for tasks requiring deeper global reasoning.
Planner: The planner orchestrates how the input is divided, what instructions each worker agent receives, and how the manager agent is prompted to unify the partial outputs. Rather than having a human specify how to break down the input, we direct the planner to do so by following steps:
  1. 1.
    Job Assignment. The planner decides how many chunks to create and which segments each worker agent will process.
  2. 2.
    Prompt Preparation. Based on the task details, the planner modifies the worker prompts so that each worker’s output can be correctly integrated downstream. It also sets up the manager prompt to combine partial results properly.
  3. 3.
    Iterative Refinement. The planner can run a brief evaluation step using some validation data to identify mispredicted cases. It then revises the prompt structure or chunking strategy to reduce errors. If repeated excessively on the same data, there is a risk of overfitting, so the planner typically does only a few refinements.
In this manner, the planner automates much of the chunking and coordination that would otherwise require manual trial and error. As we will see in the experiments, careful planner design can substantially reduce aggregator noise by ensuring each worker agent’s output is in a format that the manager agent can interpret reliably.
5 Experiments
5.1 Settings
We present experiments to assess how the three primary noise components—model noise, task noise, and aggregator noise—affect system performance under different context lengths and across multiple tasks. Section 3.4 indicates that $\phi_{\mathrm{model}}$ grows more than linearly with input length, and that synergy across subproblems ($\Phi_{\mathrm{task}}$) can impose strong requirements on any mechanism for merging partial outputs. Our experiments test the ideas on six tasks with different data scales and use several agents.
Tasks
We experiment on six diverse tasks including: Key-Value Retrieval, Math Find Number, Summarization, Dialogue Character Inference, and Open Question QA with and without choices. There tasks are based on InfiniteBench Zhang et al. (2024a) and LongBench-V2 Bai et al. (2024) but we have modified the generation and prepared different lengths of these tasks.
  • Key-Value Retrieval (KV) We randomly generated Key-Value pairs. For a given key, the task is to retrieve the value associated. This task only evaluates the capability of retrieval. The evaluation metric is accuracy. We prepare synthetic KV task to have lengths ranging from 1K to 128K.
  • Math Find Number (Math) We randomly generated a long list of integers following Gaussian distribution and the task is to find the Kth largest or smallest number in this list. Each query is slightly different in K. This task evaluates the memory and mathematical reasoning. The evaluation metric is accuracy. We prepare synthetic Math task to have lengths ranging from 1K to 128K.
  • Summarization (Sum) We provide English texts and the task is to summarize them. The ground truth summarization comes from the InfiniteBench. This task evaluates the langauge summarization. The evaluation metric is ROUGE score.
  • Open Question QA (QA-IB and QA-LB) We provide English text and some open questions for question answering. These questions are very diverse and the task requires LLM to do multi-hop reasoning to answers. The task QA-IB comes from InfiniteBench and doesn’t come with choices and the task QA-LB comes from LongBench-V2 and is provided with four choices. The evaluation metrics are respectively F1 score and accuracy.
  • Dialogue Character Inference (Char) We provide English dialogue scripts between many characters. One of the characters is masked and the task is to infer the name of the masked person. This task evaluates the capacity of language reasoning and will require LLM to focus on the interaction of people discussions to understand the relation. The evaluation metric is accuracy. This task comes from InfiniteBench.
LLM Agents
We have experimented with diverse agents choices from commercial OpenAI models and open sourced models from Meta and Alibaba. gpt-4o-2024-08-06 (gpt4o) 128K model from OpenAI. gpt-4o-mini-2024-07-18 (gpt4omini) 128K model from OpenAI. Llama-3.1-70B-Instruct (llama70b) 128K 70B LLM model from Meta Grattafiori (2024). Llama-3.2-3B-Instruct (llama3b) 128K 3B LLM model from Meta Grattafiori (2024). Qwen2.5-72B-Instruct (qwen72b) 32K 72B LLM model from Alibaba QWen Yang et al. (2024).
In the experiments where divide and conquer agents are implemented, the manager and worker agents are homogeneous and the planner agent is QWen72b. The temperature is set to 0 to minimize stochasticity during decoding.
In the following sections, we discuss respectively model noise, task noise and aggregator noise. The robustness of these observations and the practical applicability of our framework are further underscored by comprehensive supplementary analyses (detailed in Appendices D-J), which include utility of the framework, explorations of chunking variants, comparisons with RAG, evaluations across diverse model architectures, and detailed data for our main findings.
5.2 Model noise
We estimate the single-agent model noise for input $x$ as $\phi_{\mathrm{model}}(x)\;\approx\;\mathcal{D}\bigl{(}f_{\mathrm{single}}(x),\,f^{*}(x)\bigr{)},$ where $f_{\mathrm{single}}(x)$ is the output of a single model given the entire input, and $\mathcal{D}(\cdot,\cdot)$ is a distance metric such as cross-entropy or $|1-\text{accuracy}|$. By evaluating this for different input lengths $T=\|x\|$, we see how the model’s performance declines as context grows. In Figures 2(b) (Math) and 2(a) (KV), each agent processes the entire input directly (up to 128K tokens). We measure accuracy at each length; the detailed numerical scores supporting these figures are provided in Appendix D.
(a) Single agent model noise w.r.t input length: KV Task
(b) Single agent model noise w.r.t input length: Math Task
For KV retrieval (Figure 2(a)), most agents show a downward trend in accuracy as input length increases, consistent with Assumption 3.1, which posits that noise grows more than linearly. Once the length exceeds tens of thousands of tokens, performance rapidly decays. Similar observations are seen for Math (Figure 2(b)), where the loss of accuracy at longer contexts is stark for gpt4omini and llama70b. Interestingly, llama3b is already too weak at shorter lengths, so its performance saturates at near-random levels. These results reinforce our theoretical statement that $\phi_{\mathrm{model}}(x)$ can become the dominant error term for long $x$. For tasks that do not require intense synergy across the entire input, splitting the input into shorter segments (i.e., reducing $L$ per worker) should mitigate this growth, an idea formalized in Thm 3.2.
5.3 Task noise
Next, we examine how multi-agent performance depends on $\Phi_{\mathrm{task}}$ and $\Phi_{\mathrm{model}}$. We use the approximate measures:
$\Phi_{\mathrm{model}}\;\approx\;\mathcal{D}\bigl{[}f_{\mathrm{D\&C}}(x;\,\mathrm{optimal\text{-}n}),\,f_{\mathrm{D\&C}}(x;\,\mathrm{bad\text{-}n})\bigr{]},$
$\Phi_{\mathrm{task}}\;\approx\;\int_{2}^{n}\mathcal{D}\bigl{[}f_{\mathrm{D\&C}}(x;\lfloor i\rfloor),\,f_{\mathrm{single}}(x)\bigr{]}\;\mathrm{d}i.$
Here, $\Phi_{\mathrm{model}}$ is captured by the gap between choosing an effective chunk size versus a poor one, while $\Phi_{\mathrm{task}}$ is reflected by how the system output diverges from the single-agent baseline as we increase the number of chunks from 2 to $n$.
(a) KV
(b) Math
(c) QA-IB
(d) QA-LB
(e) Sum
(f) Char
Figure 3: Joint effect of task noise $\Phi_{\mathrm{task}}$ and model noise $\Phi_{\mathrm{model}}$. According to the discussion in Sec 3.4, (a) has negligible $\Phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$;(b)-(e) has dominating model noises $\Phi_{\mathrm{model}}$; (f) has dominating task noise $\Phi_{\mathrm{task}}$.
Figure 3 illustrates the performance of five agents on the 128K version of each task, varying the chunk size from 1K to 64K tokens. The comprehensive performance metrics for all tasks and models underpinning this figure are available in Appendix E. We observe three patterns that match the three regimes described in Section 3.4:
  1. 1.
    Low task noise (KV). When cross-chunk synergy is minimal, splitting yields similar results regardless of $n$. This matches the “low task noise, low model noise” regime, where $\Phi_{\mathrm{task}}$ is nearly zero and $\phi_{\mathrm{model}}$ remains small, so aggregator choices have little impact.
  2. 2.
    Dominating model noise (Math, QA, Sum). In these tasks, large input length makes single-shot usage prone to confusion, so $\phi_{\mathrm{model}}$ becomes significant. However, $\Phi_{\mathrm{task}}$ is moderate enough that a basic aggregator can handle partial results effectively. Splitting into smaller chunks thus reduces per-chunk confusion and outperforms a single-shot approach. This behavior aligns with Thm 3.2, showing that superlinear growth of $\phi_{\mathrm{model}}$ can be contained with divide-and-conquer.
  3. 3.
    Dominating task noise (Char). In this scenario, cross-chunk interactions are so extensive that partial outputs cannot capture the global context unless the aggregator reintroduces nearly the entire input. If the aggregator does not do so, performance remains low. This corresponds to the “dominating task noise” regime, where $\Phi_{\mathrm{task}}$ is large enough to overshadow the benefits of splitting.
We also explored whether introducing a small token overlap between chunks could mitigate task noise effects. However, as detailed in Appendix G, a 1K token overlap yielded mixed and generally marginal benefits for the Llama-70B model on 128K tasks, suggesting it does not fundamentally alter these noise trade-offs for the scenarios tested. Furthermore, to assess the generalizability of our noise framework, we evaluated additional diverse models across various tasks, as detailed in Appendix J. These results confirm that while effective context length (e.g., per Ruler benchmark) influences performance, the observed noise patterns and the utility of our decomposition are consistent, with task complexity significantly interacting with these factors across different architectures.
Overall, these empirical data show that the relative magnitudes of $\phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$ determine whether chunking helps or hurts. Moderate-synergy tasks benefit most from a chunk-based approach, validating our theoretical insights in Section 3.4. Apart from the high-synergy character-inference problem, divide-and-conquer methods achieve performance on par with or exceeding the strongest single-shot model, reinforcing the idea that chunking is advantageous when model noise dominates and task noise remains modest.
5.4 Aggregator noise
Finally, we examine $\Phi_{\mathrm{agg}}$, which measures how well the partial results are merged. We estimate by
$\Phi_{\mathrm{agg}}\;\approx\;\int_{2}^{n}\mathcal{D}\bigl{[}f^{\mathrm{optimal\text{-}agg}}_{\mathrm{D\&C}}(x;\lfloor i\rfloor),f^{\mathrm{bad\text{-}agg}}_{\mathrm{D\&C}}(x;\lfloor i\rfloor)\bigl{]}di,$
Here, we compare a stronger aggregator prompt (planner-based) to a weaker one (manual). Figure 4 contrasts two approaches on the Math and QA-LB tasks for large $T$: (1) a manual aggregator prompt, which simply asks each worker to solve its subproblem in isolation, and (2) a planner-based aggregator prompt, which enforces a structured approach to partial outputs. This yields a visible performance gap (shaded area in Figure 4), showing that aggregator noise can be substantially reduced via more coordinated prompts. The same effect is seen in QA-LB, where the aggregator must reconcile partial answers from each worker.
(a) Math
(b) Math
(c) QA-LB
(d) QA-LB
Figure 4: Aggregator noise across different tasks and models.
These empirical results consolidate Section 3.4 and the aggregator noise assumption (A): a robust aggregator prompt can keep $\Phi_{\mathrm{agg}}$ modest, while naive prompts inflate the final error. The planner effectively refines aggregator instructions to align worker outputs with the manager’s requirements, illustrating how a minimal overhead aggregator can preserve the theoretical advantages of chunking large inputs.
6 Conclusion
We presented a theoretical and empirical investigation of divide-and-conquer methods for large context tasks. Our framework separates the final error into task noise (cross-chunk synergy), model noise (per-chunk confusion), and aggregator noise (integration flaws). We showed that chunking can outperform single-shot usage in many tasks where superlinear noise growth appears with context length. However, if cross-chunk synergy is too large or the aggregator prompt is ineffective, chunk-based approaches may fail.
Our experiments illustrate several key behaviors central to our noise decomposition framework, including the sweet-spot trade-off between synergy (task noise) and chunk size (model noise), and the conditions under which weaker models leveraging D&C can outperform stronger single-shot models afflicted by superlinear model noise at large context lengths. We also showed that aggregator noise can be reduced by careful prompt engineering, which can be partially automated using a planner.
In future work, it would be interesting to explore more powerful aggregator models or retrieval-based approaches that reconstruct the relevant cross-chunk information. We hope this framework clarifies the interplay between context length, synergy, and multi-agent design, and provides a path toward improved performance on truly long documents.
Appendix A Justification of Assumption 3.1
Justification by elimination.
We consider different growth rates explicitly:
Sublinear growth.
If $\phi_{\mathrm{model}}(x)$ were sublinear, then the average noise per token would decrease as input length increases, which contradicts empirical observations showing performance degradation with larger contexts.
Linear growth ($O(L)$).
Suppose $\phi_{\mathrm{model}}(x)\approx cL$. Then, dividing the input of length $T$ into $n$ chunks (each of length $T/n$) would yield total model noise:
$\Phi_{\mathrm{model}}=n\cdot c\frac{T}{n}=cT,$
which exactly matches the single-shot model noise. Under strictly linear model noise, chunking would provide no direct benefit in terms of model noise alone.
However, consider the empirical evidence from a key-value (KV) retrieval task: this scenario has negligible task noise ($\Phi_{\mathrm{task}}\approx 0$) and negligible aggregation noise ($\Phi_{\mathrm{agg}}\approx 0$), as retrieval operations are completely independent across chunks. Despite this, we empirically observe a clear and substantial improvement from chunked inference.
Such empirical evidence clearly indicates that linear noise growth cannot fully explain observed performance gains. Therefore, the assumption of linear model noise must be incorrect, and the model noise must instead be strictly superlinear, consistent with polynomial growth ($O(L^{\alpha}),\alpha>1$).
Intermediate growth ($O(L\log L)$).
Suppose $\phi_{\mathrm{model}}(x)\approx cL\log L$. Dividing the input of length $L$ into $n$ chunks of length $L/n$ yields total noise:
$n\cdot c\frac{L}{n}\log\frac{L}{n}=cL(\log L-\log n).$
To examine how small the incremental gain is, consider doubling the number of chunks from $n/2$ to $n$:
$cL(\log L-\log(n/2))-cL(\log L-\log n)=cL(\log n-\log(n/2))=cL\log 2.$
Thus, each time we double the chunk number, we gain only a constant additive improvement ($cL\log 2$), which is independent of $n$. Crucially, this incremental improvement does not increase with larger $n$, but remains constant.
Empirically, however, we observe significant and increasing performance improvements as we continue to subdivide the context into smaller chunks (as shown in the KV retrieval scenario with negligible task and aggregation noises). Such empirical behavior clearly contradicts the prediction of $O(L\log L)$ complexity, which would provide diminishing returns (constant incremental improvement) upon increasing the chunk number.
Therefore, intermediate complexity growth of the form $O(L\log L)$ cannot adequately explain observed empirical benefits, indicating that the true complexity must be strictly faster (e.g., polynomial superlinear growth).
Polynomial superlinear growth ($O(L^{\alpha})$, $\alpha>1$).
Under polynomial superlinear growth, dividing input into $n$ chunks yields total noise:
$n\cdot c\left(\frac{L}{n}\right)^{\alpha}=cn^{1-\alpha}L^{\alpha}.$
Since $\alpha>1$, the factor $n^{1-\alpha}$ decreases significantly with increasing $n$, aligning closely with empirical observations of substantial improvements.
Exponential growth ($O(\exp(L))$).
Exponential complexity would quickly make inference infeasible for even moderate lengths, clearly inconsistent with the practical performance of real-world language models.
Other intermediate or exotic complexities.
While other theoretical complexities (e.g., $O(L\log^{k}L)$, $O(L^{\log L})$, $O(\exp(\sqrt{L}))$) exist, they either (1) fail to match empirical performance characteristics (like offering insufficient improvement), or (2) represent unrealistic scenarios never observed empirically in LLM inference tasks.
Thus, by explicit elimination, empirical evidence strongly supports polynomial superlinear complexity $\phi_{\mathrm{model}}(x)\geq cL^{\alpha}$ with $\alpha>1$. ∎
Appendix B Proof of Thm 3.2
Proof.
Step 1. Bounding $\Phi_{\mathrm{model}}$ via superlinear local noise reduction. Recall the definition of the aggregate model noise:
$\Phi_{\mathrm{model}}=\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}-\prod_{i=1}^{n}f_{i}^{*}(x_{i}).$
This term captures the noise introduced by the worker models $\phi_{\mathrm{model}}(x_{i})$. By Assumption 3.1, we assume the local model noise $\phi_{\mathrm{model}}(L)$ grows superlinearly with chunk length $L$, specifically $|\phi_{\mathrm{model}}(L)|\geq cL^{\alpha}$ for constants $c>0$, $\alpha>1$, and sufficiently large $L$.
When we partition the input $x$ of length $T$ into $n$ chunks $x_{i}$, each of length approximately $T/n$, the magnitude of the local noise $|\phi_{\mathrm{model}}(x_{i})|$ is significantly reduced compared to the single-shot noise $|\phi_{\mathrm{model}}(T)|\geq cT^{\alpha}$. While the exact form of $\Phi_{\mathrm{model}}$ involves combinations of these local noise terms (potentially multiplied by ideal output terms $f_{j}^{*}$), the dominant factor driving the change in the magnitude $|\Phi_{\mathrm{model}}|$ is this substantial, superlinear reduction in the individual $|\phi_{\mathrm{model}}(x_{i})|$ terms achieved through chunking.
We can characterize the scaling of $|\Phi_{\mathrm{model}}|$ by considering an upper bound proportional to the sum of the bounds on these dominant local noise contributions. Specifically, when local noises $\phi_{\mathrm{model}}(x_{i})$ are small, the first-order terms in the expansion of $\Phi_{\mathrm{model}}$ involve a sum over terms linear in each $\phi_{\mathrm{model}}(x_{i})$ (e.g., $\sum_{i}(\phi_{\mathrm{model}}(x_{i})\prod_{j\neq i}f_{j}^{*}(x_{j}))$ if $\prod$ is multiplication). The overall magnitude $|\Phi_{\mathrm{model}}|$ can thus be bounded by a term related to the sum of the local noise bounds:
$|\Phi_{\mathrm{model}}|\leq B\cdot\sum_{i=1}^{n}|\phi_{\mathrm{model}}(x_{i})|\leq B\cdot n\cdot\sup_{i}|\phi_{\mathrm{model}}(x_{i})|$
where B accounts for the influence of the $f^{*}_{j}$ terms and higher-order noise interactions (assumed reasonably bounded). Applying the local superlinear bound $|\phi_{\mathrm{model}}(L)|\leq c^{\prime}L^{\alpha}$ (using $c^{\prime}$ for an upper bound constant potentially different from $c$):
$|\Phi_{\mathrm{model}}|\leq B\cdot n\cdot c^{\prime}\left(\frac{T}{n}\right)^{\alpha}=(Bc^{\prime})\cdot n^{1-\alpha}T^{\alpha}.$
Let $C_{bound}=Bc^{\prime}$. Since $\alpha>1$, we have $01-\alpha<$ 0. Therefore, increasing the number of chunks $n$ decreases the factor $n^{1-\alpha}$, ensuring that this upper bound on $|\Phi_{\mathrm{model}}|$ grows slower than $T^{\alpha}$. Specifically, for sufficiently large $T$, we can choose $n$ such that:
$|\Phi_{\mathrm{model}}|\leq C_{bound}\,n^{1-\alpha}T^{\alpha}<cT^{\alpha}\leq|\phi_{\mathrm{model}}(T)|.$
Thus, the total model noise introduced by the D&C approach, $|\Phi_{\mathrm{model}}|$, can be made strictly smaller than the noise magnitude of the single-shot approach, $|\phi_{\mathrm{model}}(T)|$, by appropriately choosing the number of chunks $n$.
Step 2. Bounded mismatch $\Delta^{*}(x)$ (Assumption A2). We explicitly assume there exists a constant $\Delta_{0}$ such that:
$|\Delta^{*}(x)|=\left|\prod_{i=1}^{n}f_{i}^{*}(x_{i})-\prod_{i=1}^{n}g^{*}(x_{i})\right|\leq\Delta_{0},$
independent of the input length $T$ and chunking strategy. Thus, $\Delta^{*}(x)$ does not scale with input size.
Step 3. Bounded aggregator noise $\Phi_{\mathrm{agg}}$ (Assumption A3). We explicitly assume there exist constants $M_{\mathrm{agg}}>0$ and $n_{\mathrm{agg}}>0$ such that:
$|\Phi_{\mathrm{agg}}|\leq M_{\mathrm{agg}},\quad\text{for }n\leq n_{\mathrm{agg}}.$
This ensures aggregation noise remains bounded for a controlled number of chunks.
Step 4. Bounded cross-chunk mismatch $\Phi_{\mathrm{task}}$ (Assumption A4). We explicitly assume there exist constants $M_{\mathrm{task}}>0$ and $n_{\mathrm{task}}>0$ such that:
$|\Phi_{\mathrm{task}}|=|\phi_{\mathrm{task}}(x_{1},\dots,x_{n})|\leq M_{\mathrm{task}},\quad\text{for }n\leq n_{\mathrm{task}}.$
This ensures task noise does not grow significantly when using sufficiently few chunks.
Step 5. Combining all terms. By combining Steps 1–4, the total error of the D&C approach is:
$|f_{\mathrm{D\&C}}(x;n)-g(x)|$ $\leq|\Delta^{*}(x)|+|\Phi_{\mathrm{model}}|+|\Phi_{\mathrm{agg}}|+|\Phi_{\mathrm{task}}|$
$\leq\Delta_{0}+|\Phi_{\mathrm{model}}|+M_{\mathrm{agg}}+M_{\mathrm{task}}.$
Using the bound derived in Step 1, for $n$ chosen appropriately (e.g., $n\approx T/L_{0}$, subject to $n\leq\min(n_{\mathrm{agg}},n_{\mathrm{task}})$):
$|f_{\mathrm{D\&C}}(x;n)-g(x)|\leq\Delta_{0}+(C_{bound}\,n^{1-\alpha}T^{\alpha})+M_{\mathrm{agg}}+M_{\mathrm{task}}.$
Since $\alpha>1$, the term $C_{bound}\,n^{1-\alpha}T^{\alpha}$ grows slower than $T^{\alpha}$. The terms $\Delta_{0},M_{\mathrm{agg}},M_{\mathrm{task}}$ are constants. The single-shot noise magnitude $|\phi_{\mathrm{model}}(T)|$ grows at least as fast as $cT^{\alpha}$. Therefore, for sufficiently large $T$, the sum of the bounded terms and the slower-growing model noise term will be less than the rapidly growing single-shot noise:
$\Delta_{0}+(C_{bound}\,n^{1-\alpha}T^{\alpha})+M_{\mathrm{agg}}+M_{\mathrm{task}}<cT^{\alpha}\leq|\phi_{\mathrm{model}}(T)|.$
Thus,
$|f_{\mathrm{D\&C}}(x;n)-g(x)|<|\phi_{\mathrm{model}}(T)|,$
demonstrating that the divide-and-conquer inference error can be made smaller than the single-shot model noise magnitude for sufficiently large $T$. ∎
Appendix C Sample prompts for planner, manager and worker agents
The planner provides an automated way to transform a raw prompt into structured instructions for both worker and manager agents. It parses the original task prompt, identifies key requirements, and then generates appropriate prompts for the workers and the manager. This reduces manual overhead in designing separate prompts, since the planner consistently creates focused directives for each agent.
Planner Prompts Here is a prompt that I use for one LLM agent. Now since my input is too long, I divide it into multiple chunks and feed one chunk to one worker agent. Suppose now that you are the manager agent to synergize the responses of multiple worker agents in order to complete the prompt task. Write the prompt for each worker agent and the prompt for manager agent to synthesize the answer.
{RAW TASK PROMPT}
C.1 Example: Summerization
To further illustrate the planner, we use summerization task as another example. We first show the manual (default) prompts for worker and manager agents as below.
Raw prompt from benchmark Read the book paragraphs below and summarize. Book: {context}
Worker agent prompt (manual) Read the book paragraphs below and summarize. Book: {text_chunk}
Manager agent prompt (manual) There is a list of candidates. Each candidate has been given one sequential part of a book and summarize their own parts. You have been provided with their responses. Summarize their summaries into one. Responses from candidates: {response_from_workers}
Now we use the designed planner to automatically prepare the appropriate prompts for workers and manager. We feed the raw prompt from benchmark to the planner.
Worker agent prompt (planner) You are part of a team tasked with summarizing a portion of a book. Below is an excerpt from the book assigned to you. Please provide a concise summary of this specific section, highlighting key points and themes. Excerpt: {book_chunk} Please ensure your summary is clear and focuses on the main ideas presented in the excerpt.
Manager agent prompt (planner) You are the manager agent responsible for synthesizing the summaries provided by multiple worker agents. Each worker agent has been given a different section of the book to summarize. Your task is to combine these individual summaries into a coherent and comprehensive summary of the entire book section provided to the worker agents. Responses from Worker Agents: {response_from_workers} Please integrate these summaries, ensuring that the overall summary captures the main points, themes, and any important details necessary for understanding the book’s content as a whole.
C.2 Example: QA
Raw prompt from benchmark {context} How do users most likely plan to dress for a master’s graduation? A. Wear a tuxedo or robe. B. Dress business casual. C. Wear business attire. D. Wear whatever clothes they feel comfortable in.
Worker agent prompt (planner) {context} You are a worker agent. You will receive a portion of data related to how users plan to dress for a master’s graduation. Your task is to analyze the provided information and offer insights based on your assigned chunk. Focus on one of the possible dress options and any relevant context or rationale behind the choice. The options include: A. Wear a tuxedo or robe, B. Dress business casual, C. Wear business attire, D. Wear whatever clothes they feel comfortable in. Please focus only on any details in your data relevant to one specific option and clearly summarize its findings and rationale. Provide examples if necessary.
Manager agent prompt (planner) You are the Manager Agent. You will receive responses compiled from several Worker Agents, each analyzing different portions of input data related to dressing for a master’s graduation. Your task is to review and synthesize these individual analyses into a comprehensive overview. Consider each worker agent’s insights on the various dress options: A. Tuxedo or robe, B. Business casual, C. Business attire, D. Comfortable clothes. Highlight the most popular trends and any notable reasons supporting these choices. Pay attention to diversity in preferences, cultural considerations, and any statistical or anecdotal evidence provided by worker agents. Your final output should deliver a clear synthesis on how users most likely plan to dress for this occasion, potentially indicating preferences, influences, and recommendations. Responses from candidates: {response_from_workers}
C.3 Example: Math
Raw prompt (Math) {context} (long list of numbers) What is the 2nd smallest number?
Worker agent prompt (planner) You are a worker agent tasked with analyzing a portion of a list of numbers. From the chunk of numbers provided below, identify and return the two smallest numbers in ascending order. Here is your list: {text_chunk} Please output the two smallest numbers you find in your assigned list.
Manager agent prompt (planner) You are the manager agent tasked with synthesizing results from multiple worker agents. Each worker agent has provided the two smallest numbers from their respective chunks of a larger list. Your task is to compile these results and determine the 2nd smallest number from the combined outputs of all worker agents. Responses from worker agents: {response_from_workers} Please analyze these results and provide the 2nd smallest number from the original list.
Appendix D Model noise results
This section provides the detailed numerical data that form the basis for Figure 2 presented in Section 5.2 of the main paper. Figure 2 illustrates the performance degradation of single-agent Large Language Models (LLMs) as the input context length increases, a phenomenon we attribute to model noise ($\phi_{model}(x)$). Table 1 tabulates the specific accuracy scores for the evaluated models (gpt4o, gpt4omini, llama3b, and llama70b) on the Key-Value Retrieval (KV) and Math Find Number (Math) tasks, across a range of input lengths from 1K to 128K tokens. These raw scores directly support the visualizations and analysis concerning the growth of model noise with increasing context size discussed in the main text.
Table 1: Model noise results. Scores for both tasks are accuracy.
Task Task length gpt4o gpt4omini llama3b llama70b
KV 1K 1.00 1.00 1.00 1.00
KV 2K 1.00 1.00 0.98 1.00
KV 4K 1.00 1.00 0.95 1.00
KV 8K 1.00 1.00 0.88 1.00
KV 16K 1.00 1.00 0.66 1.00
KV 32K 1.00 0.99 0.24 1.00
KV 64K 1.00 0.86 0.01 0.91
KV 128K 1.00 0.60 0.01 0.15
Math 1K 0.67 0.71 0.23 0.63
Math 2K 0.68 0.65 0.19 0.65
Math 4K 0.71 0.64 0.19 0.66
Math 8K 0.69 0.61 0.10 0.55
Math 16K 0.62 0.54 0.06 0.55
Math 32K 0.63 0.51 0.06 0.52
Math 64K 0.54 0.37 0.04 0.39
Math 128K 0.33 0.11 0.00 0.09
Appendix E Task noise results
This section contains the detailed numerical results corresponding to Figure 3 in Section 5.3 of the main paper. Figure 3 illustrates the joint effect of task noise ($\Phi_{task}$) and model noise ($\Phi_{model}$), showcasing how performance varies for different LLM agents on 128K-token length tasks when the effective model context length (i.e., chunk size in the D&C setting) is altered. Table 2 lists the precise performance metrics (accuracy, F1 score, or ROUGE score, depending on the task) for the models gpt4omini, llama70b, and qwen72b. These scores are provided for all six benchmark tasks (KV, Math, QA-IB, QA-LB, Sum, Char) across various model context lengths ranging from 1K to 64K tokens (when processing a total input of 128K tokens). This data underpins the analysis of the three noise regimes and the conditions under which D&C approaches are advantageous, as discussed in the main text.
Table 2: Task noise results. Scores for all tasks are accuracy except sum which is ROUGE and qaib which is F1 score.
Model Context Char KV Math QA-IB QA-LB Sum
gpt4omini 1000 0.12 0.99 0.49 0.22 0.33 0.08
gpt4omini 2000 0.12 0.99 0.50 0.27 0.44 0.11
gpt4omini 4000 0.15 0.99 0.55 0.39 0.44 0.15
gpt4omini 8000 0.17 0.99 0.54 0.36 0.48 0.14
gpt4omini 16000 0.15 1.00 0.50 0.37 0.44 0.15
gpt4omini 32000 0.18 1.00 0.55 0.42 0.46 0.11
gpt4omini 64000 0.16 0.98 0.45 0.39 0.38 0.11
llama70b 1000 0.04 0.99 0.47 0.44 0.15 0.16
llama70b 2000 0.05 0.96 0.49 0.55 0.36 0.19
llama70b 4000 0.05 0.99 0.50 0.55 0.46 0.23
llama70b 8000 0.07 0.98 0.50 0.56 0.33 0.28
llama70b 16000 0.07 1.00 0.57 0.63 0.26 0.24
llama70b 32000 0.13 1.00 0.53 0.54 0.28 0.23
llama70b 64000 0.17 0.91 0.39 0.41 0.31 0.21
qwen72b 1000 0.08 0.98 0.33 0.32 0.31 0.08
qwen72b 2000 0.07 0.95 0.46 0.40 0.28 0.18
qwen72b 4000 0.07 0.97 0.41 0.44 0.41 0.29
qwen72b 8000 0.15 0.98 0.40 0.48 0.54 0.23
qwen72b 16000 0.15 1.00 0.36 0.48 0.46 0.18
qwen72b 32000 0.13 0.88 0.31 0.42 0.46 0.19
Appendix F Framework utility of predictive parameter estimation
A important measure of a theoretical framework’s value lies in its empirical utility, extending beyond explanatory power to offer practical benefits. This section discusses both the explanatory strengths of our noise decomposition framework and its predictive utility in efficiently determining optimal parameters for Divide and Conquer (D&C) strategies.
F.1 Explanatory Power
Our framework provides a principled understanding of why and when D&C methods succeed or fail in the context of long-document processing by Large Language Models (LLMs). By decomposing the overall error into distinct model noise ($\Phi_{model}$), task noise ($\Phi_{task}$), and aggregator noise ($\Phi_{agg}$), the framework facilitates the disentanglement of various contributing factors to performance. This structured understanding has been instrumental in deriving the key insights presented throughout this paper, such as the conditions favoring chunking and the impact of superlinear model noise growth.
F.2 Predictive Utility for Optimal Chunk Size Estimation
A practical challenge in deploying D&C methods is the selection of an optimal chunk size, which often requires costly exhaustive grid searches over various configurations. We conducted new experiments to evaluate whether our framework’s insights could guide a more efficient estimation of this parameter, particularly in regimes dominated by model noise (Case 3 in Section 3.4), where chunk size significantly impacts performance. In other regimes (Case 1 and Case 2), performance tends to be less sensitive to chunk size variations, making this estimation less critical.
We focused on tasks where model noise is a dominant factor (QA-IB and Summarization, as identified in Figure 3) and attempted to estimate the optimal chunk size by evaluating performance on a very small number of randomly selected document samples (3, 5, or 10 samples) for each potential chunk size configuration. The D&C performance achieved with the chunk size selected via this minimal sampling approach was then compared to the optimal performance found through an exhaustive grid search over all samples for all chunk sizes.
The results for the QA-IB and Summarization (Sum) tasks are presented in Table 3 and Table 4, respectively. These results demonstrate that even with minimal sampling (as few as 3-5 samples per configuration), it is possible to identify chunk sizes that yield near-optimal D&C performance. This suggests a pathway to significant computational savings when setting up D&C pipelines for new tasks or models, by avoiding exhaustive searches. For example, in the QA-IB task with ‘llama70b‘, using just 5 samples per chunk size configuration allowed us to identify a 16K chunk size yielding a score of 0.63, identical to the optimal score and chunk size found via exhaustive search. Similarly, for ‘qwen72b‘ on the Summarization task, 5 samples were sufficient to find the optimal 4K chunk size and achieve the optimal score of 0.29.
Table 3: Predictive Utility on QA-IB: Optimal Chunk Size Estimation (Score (Optimal Chunk Size Found)). Total task length 128K tokens.
Model 3-sample 5-sample 10-sample Optimal after Exhaustive Search
gpt4omini 0.38 (64K) 0.42 (32K) 0.42 (32K) 0.42 (32K)
llama70b 0.55 (2K) 0.63 (16K) 0.63 (16K) 0.63 (16K)
qwen72b 0.40 (2K) 0.48 (16K) 0.48 (8K) 0.48 (8K & 16K)
Table 4: Predictive Utility on Summarization (Sum): Optimal Chunk Size Estimation (Score (Optimal Chunk Size Found)). Total task length 128K tokens.
Model 3-sample 5-sample 10-sample Optimal after Exhaustive Search
gpt4omini 0.15 (16K) 0.14 (8K) 0.14 (8K) 0.15 (4K & 16K)
llama70b 0.23 (32K) 0.24 (16K) 0.28 (8K) 0.28 (8K)
qwen72b 0.23 (8K) 0.29 (4K) 0.29 (4K) 0.29 (4K)
Rationale for Efficient Estimation
The feasibility of accurately estimating optimal chunk sizes from sparse samples can be understood through our framework’s characterization of model noise. We posit that model noise ($\phi_{model}(x)$) exhibits superlinear growth with context length $L=||x||$. It is reasonable to infer that this underlying noise function $\phi(L)$ is, in practice, often (or can be approximated as) monotonic with respect to $L$ over relevant ranges. Consequently, within the D&C paradigm, reducing the per-worker chunk size is expected to lead to a predictably monotonic (or near-monotonic) decrease in the dominant accumulated model noise component, $\Phi_{model}$.
When $\Phi_{model}$ dominates (and task noise $\Phi_{task}$ is relatively small and aggregator noise $\Phi_{agg}$ is managed), the overall D&C system error as a function of chunk size is likely to exhibit a discernible optimal region (e.g., a convex-like shape or a region where decreasing model noise is balanced by slowly increasing task/aggregator noise) rather than fluctuating randomly and unpredictably. Therefore, a small number of samples across different chunk sizes can provide sufficient information to sketch the rough contour of this error curve, allowing for the identification of the approximate location of this optimal trade-off point without exhaustive evaluation.
Appendix G The impact of overlap in chunking strategies
In our main analysis, we primarily considered non-overlapping chunks. To further explore variations in chunking methodologies, we investigated the impact of introducing a modest overlap between adjacent chunks. This strategy is sometimes employed with the aim of mitigating potential information loss at chunk boundaries, which could otherwise contribute to task noise ($\Phi_{task}$). We conducted experiments using the Llama-70B model on several 128K-token long-context tasks, comparing non-overlapping chunks with chunks having a 1K token overlap (base chunk sizes were 4K and 16K tokens before overlap).
The results are presented in Table 5.
Table 5: Impact of 1K Token Overlap on Llama-70B Performance (128K Tasks)
Llama-70B Strategy KV QA-IB Sum Char
16K chunks, no overlap 1.00 0.63 0.24 0.07
16K chunks, 1K overlap 1.00 0.54 0.25 0.07
4K chunks, no overlap 0.99 0.55 0.23 0.05
4K chunks, 1K overlap 1.00 0.53 0.19 0.05
The inclusion of a 1K token overlap led to mixed and generally marginal changes in performance. While slight improvements were noted in some specific instances, other cases showed minor degradations or no discernible effect. This suggests that while a small overlap might offer a limited benefit for certain local cross-chunk dependencies, it does not consistently or substantially alter the overall performance trade-offs identified by our framework for these tasks. Moreover, extensive overlap could introduce processing redundancies or even conflicting information for the aggregator, potentially increasing aggregator noise ($\Phi_{agg}$). For the configurations tested, the benefits of this overlap strategy did not appear to consistently outweigh the increased complexity or computational cost.
Appendix H Comparative Analysis with Retrieval-Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) represents a prominent alternative for addressing long-context tasks by first retrieving presumptively relevant segments of the input, which are then processed by the LLM. To situate our Divide and Conquer (D&C) framework in relation to this approach, we performed a comparative analysis. We applied RAG to the single-shot baseline models for the 128K-token versions of our benchmark tasks, employing both BM25 (sparse retrieval) and ‘all-mpnet-base-v2‘ embeddings (dense retrieval). Based on the task query, the RAG pipeline retrieved document sections that were then provided as context to the LLM.
The performance of RAG is detailed in Table 6.
Table 6: Comparison of D&C (Implicit via "noRAG" baseline) and RAG Strategies
Model Method KV QA-IB QA-LB Sum Char
gpt4omini noRAG (Baseline) 0.60 0.23 0.31 0.13 0.19
RAG (BM25) 0.79 0.13 0.23 0.12 0.11
RAG (mpnet) 0.73 0.19 0.27 0.13 0.13
llama70b noRAG (Baseline) 0.15 0.56 0.23 0.19 0.18
RAG (BM25) 0.81 0.14 0.23 0.15 0.10
RAG (mpnet) 0.77 0.38 0.26 0.15 0.14
For several tasks, RAG was notably less effective than both our D&C methodology and the original single-shot baseline. Tasks such as summarization or character inference, which often require a global understanding or synthesis of diffuse information not easily targeted by a simple query, proved challenging for the retrieval step. Inaccurate or incomplete retrieval frequently provided the LLM with a partial or skewed view of the overall context, leading to degraded performance. These findings highlight the sensitivity of RAG to retrieval quality and its applicability to tasks where key information is sparse or not easily queryable. Our D&C approach, by processing the entirety of the text through structured decomposition, can offer greater robustness for such scenarios.
Appendix I Relation to Architectural Context Extension Methods
Alongside algorithmic approaches like D&C, significant research efforts focus on extending the native context window capabilities of LLM architectures, often through modifications to positional encodings (e.g., RoPE scaling) or continued training on longer sequences. It is noteworthy that one of our primary baseline models, QWen2.5-72B-Instruct, already incorporates such advanced context extension techniques and has undergone training with them.
To further understand the landscape, particularly regarding training-free extension methods, we examined the behavior of Llama-2 (a model with a nominal 4K token context window) when its context was extended to 32K tokens via RoPE scaling, without additional fine-tuning. The results are shown in Table 7.
Table 7: Performance of Llama-2 with RoPE Scaled Context Extension (Training-Free)
Model Tested CTX KV QA-IB Char
Llama2-4K 4K 0.47 0.23 0.00
Llama2-32K (Scaled) 8K 0.42 0.19 0.00
16K 0.23 0.17 0.03
32K 0.12 0.13 0.00
The data indicate a decline in performance as the input length significantly surpassed the model’s original training parameters. This aligns with broader observations in the field (e.g., the LongRoPE study) suggesting that while training-free scaling offers some benefit, achieving robust performance at substantially extended context lengths typically requires further training or architectural adaptation. The D&C framework presented in this paper is complementary, providing a strategy to handle extensive sequences that can be applied irrespective of a model’s inherent maximum context length.
Appendix J Connection to effective context length RULER benchmarks
The efficacy of any long-context strategy is intrinsically linked to the underlying LLM’s capabilities. Benchmarks like Ruler aim to quantify the "effective context length" of models, often focusing on retrieval-based tasks. To assess the generalizability of our noise decomposition framework across a spectrum of models and task types beyond simple retrieval, we conducted additional evaluations. We selected three models with differing architectural characteristics and reported long-context capabilities: Mistral-7B-Instruct-v0.2, Qwen1.5-72B-Chat, and GPT-4-Turbo.
These models were evaluated on several of our benchmark tasks with varying input context lengths. The performance patterns are summarized in Table 8.
Table 8: Performance of Diverse Models with Varying Context Lengths
Model Ruler CTX (Approx.) Tested CTX KV QA-LB Sum
Mistral-7B 16K 4K 0.03 0.230 0.090
8K 0.00 0.290 0.090
16K 0.11 0.170 0.090
Qwen1.5-72B-Chat 32K 4K 0.92 0.360 0.110
8K 0.97 0.260 0.130
16K 0.95 0.310 0.110
32K 0.93 0.270 0.150
GPT-4 (Turbo) 64K+ 4K 1.00 0.431 0.088
8K 1.00 0.461 0.091
16K 1.00 0.435 0.092
32K 0.98 0.399 0.085
64K 0.87 0.384 0.081
128K 0.77 0.373 0.074
While a longer effective context is generally advantageous, the specific manifestation of model noise ($\Phi_{model}$) and task noise ($\Phi_{task}$) remains highly task-dependent. Complex reasoning or generation tasks can stress long-context capabilities differently than retrieval tasks. These observations reinforce the utility of our noise decomposition framework as an analytical tool. The framework helps explain performance variations by considering the interplay of the task’s intrinsic decomposability, the model’s confusion threshold with increasing context, and the aggregator’s ability to synthesize information, providing insights that extend beyond a single metric of effective context length.
Appendix K Framework implementation in more details
We present a simple implementation of the framework as in Figure 1. Our implementation consists of three parts: a planner for automated query planning, numerous worker agents, and single manager agent. For a long context task, we divide it into chunks of approximately equal length and feed each chunk to one worker agent. Thus, we need to design how worker agents process their chunks and how manager agent could synthesize the appropriate answer from the interaction.
K.1 Worker Agent
In our framework, the long context sequence is divided into multiple segments, with each worker agent independently responsible for processing one specific segment. This design isolates the processing of individual segments, deliberately omitting any consideration of dependencies across segments. The goal is to simplify each worker’s task by restricting its scope to a single context segment. While our implementation employs homogeneous worker agents for simplicity, this setup can be extended to accommodate heterogeneous agents with varying capacities in more complex scenarios.
K.2 Manager Agent
The manager agent is tasked with aggregating the outputs generated by the worker agents. It analyzes the responses received and synthesizes a unified, refined output. For simplicity in our current design, we use the same model architecture for the manager agent as for the worker agents. However, this design can be adapted to allow for specialized managerial processing in more advanced applications.
K.3 Planner
The planner plays a central role in the coordination of the system. While we have outlined the basic structure of worker and manager agents, the specific responsibilities of each agent will vary depending on the task and the data. Instead of manually assigning tasks to the agents, we introduce a planner that autonomously determines the optimal distribution of jobs between the worker and manager agents. The planner is given the original task description (e.g., "identify the second largest number in this list") and is responsible for generating appropriate prompts for each agent, taking into account the segmentation of the long sequence. Moreover, human-designed agent workflows typically involve a trial-and-error approach: initial job prompts are tested on a small dataset, the failures are analyzed, and the design is iteratively refined. We integrate this iterative process into the planner by having it perform an initial zero-shot job assignment, followed by an evaluation on holdout validation data. The planner receives feedback on mispredicted cases and revises the job prompts accordingly. While this iterative process can be repeated to improve performance, there is a risk of overfitting if carried out for too long. In our experiments, we iterate only once to avoid trivial problems.